HIGH LEVEL AB INITIO INVESTIGATION OF INTERCOMBINATION BANDS OF C_{2}

\underline{DAMIAN L. KOKKIN}, GEORGE B. BACSKAY and TIMOTHY W. SCHMIDT, School of Chemistry, University of Sydney, NSW 2006, Australia.

High level ab initio calculations were carried out on the X^{1}Σ_{g}^{+}, A^{1}Π_{u}, a^{3}Π_{u}, b^{3}Σ_{g}^{−}, c^{3}Σ_{u}^{+} and d^{3}Π_{g} levels of the C_{2} radical. Calculations were carried out at the MRCI level with full-valence CASSCF reference wavefunctions and basis sets up to aug-cc-pv6z. The resulting wavefunctions and potential energy surfaces were employed to calculate Einstein coefficients for allowed transitions and intercombination bands. These values were used to model the behaviour of the C_{2} system in interstellar and circumstellar environments, paying particular attention to the Red Rectangle (HD 44179).